# Identifiability of nonlinear ODE Models with Time-Varying Parameters: the General Analytical Solution and Applications in Viral Dynamics

### v4

@@ 10.4 Fourth contribution
We proceed as in the proof of Theorem 4.2. We need to prove that the vector field:
$\xi^{X}:=\left[\begin{array}[]{c}\xi\\ {}^{u}\chi\\ \end{array}\right],$ (B.1)
with ${}^{u}\chi$ defined in (4.6), belongs to the null space of $\mathcal{O}^{X}$ in (A.3).
Let us consider the first $n_{o}$ generators of $\mathcal{O}^{X}$. As $\xi\in\mathcal{O}^{T}$, we have:
$\frac{\partial\theta_{i}}{\partial x}\cdot\xi=0,~{}~{}~{}i=1,\ldots,n_{o}.$
Hence,
$\nabla\theta_{i}\cdot\xi^{X}=\frac{\partial\theta_{i}}{\partial x}\cdot\xi+\frac{\partial\theta_{i}}{\partial w}\cdot~{}^{u}\chi=$
$0+\underbrace{\left[0,\ldots,0\right]}_{m_{w}}\cdot~{}^{u}\chi=0,~{}~{}~{}i=1,\ldots,n_{o}.$
as $\theta_{i}$ is independent of $w$ (it only depends on $x$).
Let us consider the remaining $m$ generators. We have:
.$\nabla\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}\widetilde{h}_{l}\cdot\xi^{X}=\frac{\partial\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}\widetilde{h}_{l}}{\partial x}\cdot\xi+\sum_{k=1}^{m}\frac{\partial\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}\widetilde{h}_{l}}{\partial w_{k}}~{}^{u}\chi_{k}$ (B.2)
On the other hand,
.$\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}\widetilde{h}_{l}=\accentset{\mbox{\large.}}{\mathcal{L}}_{g^{0}}\widetilde{h}_{l}+\sum_{j=1}^{m}(\mathcal{L}_{g^{j}}\widetilde{h}_{l})w_{j},~{}~{}~{}l=1,\ldots,m$ (B.3)
By using (B.3) in (B.2) and by using (4.7) we obtain:
.$\nabla\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}\widetilde{h}_{l}\cdot\xi^{X}=\xi^{0}_{l}+\sum_{j=1}^{m}\xi^{j}_{l}w_{j}+\sum_{k=1}^{m}\mathcal{L}_{g^{k}}\widetilde{h}_{l}~{}^{u}\chi_{k}=$
$\xi^{0}_{l}+\sum_{j=1}^{m}\xi^{j}_{l}w_{j}+\sum_{k=1}^{m}~{}^{m}\mu^{k}_{l}~{}^{u}\chi_{k}.$
Now, by using the expression of ${}^{u}\chi_{k}$ in (4.6) and that ${}^{m}\mu$ is the inverse of ${}^{m}\nu$ (i.e., $\sum_{k=1}^{m}~{}^{m}\mu^{k}_{l}~{}^{m}\nu^{i}_{k}=\delta^{i}_{l}$), we finally obtain:
.$\nabla\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}\widetilde{h}_{l}\cdot\xi^{X}=0.$

### v5

@@ 9.4 Fourth contribution
From (3.9) we have:
$w_{k}=\sum_{\beta=0}^{m}\nu_{k}^{\beta}\left(v_{\beta}-\sum_{l=m+1}^{m_{w}}w_{l}~{}\mathcal{L}_{g^{l}}\widetilde{h}_{\beta}\right).$
By using the condition in (4.4), we obtain:
$w_{k}=\sum_{\beta=0}^{m}\nu_{k}^{\beta}v_{\beta},$
and:
$v_{\beta}=\sum_{\alpha=0}^{m}\mu^{\alpha}_{\beta}w_{\alpha}.$
Hence:
${}^{u}\chi_{k}=\sum_{\beta=0}^{m}\mathcal{L}_{\xi}\nu_{k}^{\beta}v_{\beta}=\sum_{\beta=0}^{m}\mathcal{L}_{\xi}\nu_{k}^{\beta}\sum_{\gamma=0}^{m}\mu^{\gamma}_{\beta}~{}w_{\gamma}=-\sum_{\beta=0}^{m}\nu_{k}^{\beta}\sum_{\gamma=0}^{m}\left(\mathcal{L}_{\xi}\mu^{\gamma}_{\beta}\right)w_{\gamma}=$
$-\sum_{\beta=0}^{m}\nu_{k}^{\beta}\left(\sum_{\gamma=0}^{m}\left(\mathcal{L}_{\xi}\mathcal{L}_{g^{\gamma}}\widetilde{h}_{\beta}\right)w_{\gamma}\right)=-\sum_{i=1}^{m}\nu_{k}^{i}\left(\left(\mathcal{L}_{\xi}\mathcal{L}_{g^{0}}\widetilde{h}_{i}\right)+\sum_{j=1}^{m}\left(\mathcal{L}_{\xi}\mathcal{L}_{g^{j}}\widetilde{h}_{i}\right)w_{j}\right)=$
$-\sum_{i=1}^{m}\nu_{k}^{i}\left(\xi^{0}_{i}+\sum_{j=1}^{m}\xi^{j}_{i}~{}w_{j}\right).$